OpenStudy (anonymous):

How do I solve this inequality? 0.25r-0.125 + 0.5r=0.5 + r

OpenStudy (anonymous):

Have you tried solving this yet?

OpenStudy (anonymous):

Yes I have gone through it like 5 times and I always end up with something weird like 0.125=0.25r

OpenStudy (anonymous):

So I did this a few times. I got to the answer -2.5.

OpenStudy (anonymous):

How did you wind up getting that? After I add all my like terms I get lost. Thanks again guys for helping me out I really appreciate it.

OpenStudy (anonymous):

Heres what I did: 0.25r - 0.125 + 0.5r = 0.5 + r I combined like numbers then add: 0.25r + 0.5r - 0.125 = 0.5 + r 0.75r - 0.125 = 0.5 +r I then subtracted r from the right side to the left. 0.75r - r -0.125 = 0.5 -0.25r - 0.125 = 0.5 I added -0.125 to the right [ -0.125 + 0.5 = 0.625 ] -0.25r = 0.625 I divided to get r alone and for the value of r 0.625 / -0.25 = -2.5 r = -2.5 To check, plug -2.5 into all areas of r to get the same answer of -2 on both sides.

OpenStudy (anonymous):

Wow. Thank you for taking the time to check these. I might be awhile checking all this but thanks rbeck & isaiah. You guys helped alot

OpenStudy (anonymous):

No problem. Just remember to always check your answer. Plug the value into the problem. You should get the same answer on both sides. Also remember that .... theres 0's that is literally not showing because they are invisible and in MATH ... 0 always go blank! Haha. (That just make things very confusing.) Glad I was able to help ! Hi^5.

OpenStudy (anonymous):

Back at you man! Thanks